OpenStudy (anonymous):

square root 5n-1-6=-4

OpenStudy (anonymous):

it's not clear

OpenStudy (anonymous):

\[\sqrt{5n-1-6}=-4\]?

OpenStudy (anonymous):

plz draw it

OpenStudy (anonymous):

yes, \[\sqrt{5n-1-6} =-4\]

OpenStudy (anonymous):

what is -1-6?

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

okay. let me write this down.

OpenStudy (anonymous):

yes. so what the next step? Thank you.

OpenStudy (anonymous):

\[\sqrt{5n-7}=-4\]we can't simplify the inside of the square root any further so the next step is squaring both sides.

OpenStudy (anonymous):

This will get rid of the square root

OpenStudy (anonymous):

5x-7=16 5x=-1 x=-1/5

OpenStudy (anonymous):

DUDE!!!! why did you give the answer @moha_10

OpenStudy (anonymous):

what should i do ?? i'm just want to make sure what i did

OpenStudy (anonymous):

@moha_10 It's not about you. We can care less if you know how to solve we want to the asker to be able to solve it.

OpenStudy (anonymous):

After squaring both sides, what's the next step? Please and Thank you.

OpenStudy (anonymous):

we just transfer the numbers to another side

OpenStudy (anonymous):

like \[5n-7=16\]

OpenStudy (anonymous):

great @kitsune0724 Now we have to add 7 to both sides. This will get rid of the -7 from the left hand side.

OpenStudy (anonymous):

What do we get now?

OpenStudy (anonymous):

23

OpenStudy (anonymous):

\[5n=23\]great now we want to get rid of that 5. The way to do this is by dividing both sides by 5

OpenStudy (anonymous):

What do we get now?

OpenStudy (anonymous):

4.6?

OpenStudy (anonymous):

\[n=4.6\]Yes good job!

OpenStudy (anonymous):

thank you so much!